OpenStudy (anonymous):

find the least value of 4(cosec)2@+9(sin)2@ step by step please here 2 is square

OpenStudy (anonymous):

1. This is totally in the wrong section dude :-D You posted this in Geography 2. Did you mean this? \(4 \csc^2\theta + 9 \sin^2\theta \) 3. Step-by-step: \(\large\frac{d}{d\theta }(4 \csc^2\theta + 9 \sin^2\theta )\) \(9\large (\frac{d}{d\theta }(\sin^2\theta ))+4 (\frac{d}{d\theta }(\csc^2\theta ))\) \(9\large (\frac{d}{d\theta }(\sin^2\theta ))+4 (2 \csc\theta (\frac{d}{d\theta }(\csc\theta )))\) \(9\large (2 \sin\theta (\frac{d}{d\theta }(\sin\theta )))+8 \csc(x) (\frac{d}{d\theta }(\csc\theta ))\) \(\large18 \sin\theta (\frac{d}{d\theta }(\sin\theta ))+8 \csc\theta (-\cot\theta \csc\theta )\) \(\large18 \sin\theta \cos\theta -8 \cot\theta \csc^2\theta \) Questions?